A Complete Bibliography of Publications in Reviews of Modern Physics (1920–1929)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

02 November 2023 Version 1.03

Title word cross-reference

Constants [Bir29b, Bir29a]. Corpuscular [Com29].

Electricity [Dar29]. Errata [Bir29a, Kem30].

General [Bir29a, Bir29b, Kem29, Kem30].

Ionization [Moh29].

Light [Com29].

Matter [Dar29]. Mechanics [Kem29, Kem30].

Part [Kem29, Kem30]. Photo [Moh29]. Photo-Ionization [Moh29]. Physical [Bir29a, Bir29b]. Principles [Kem29, Kem30]. Probable [Bir29b, Bir29a]. Properties [Com29].

Quantum [Kem29, Kem30].

Radiation [Dar29]. Rays [Osg29]. Recombination [Moh29].

Soft [Osg29]. Spectroscopy [Osg29]. Statistical [Dar29].

Theories [Dar29].

Values [Bir29a, Bir29b].

X [Osg29]. X-Rays [Osg29].

References

Birge:1929:EPV

[Bir29a] Raymond T. Birge. Errata: Probable Values of General Physical Constants. Reviews of Modern Physics, 1(2):241, October 1929. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
1.241;   http://rmp.aps.org/abstract/RMP/v1/i2/p241_1. See

[Bir29b].

Birge:1929:PVG

[Bir29b] Raymond T. Birge. Probable values of the general physical constants. Reviews of Modern Physics, 1(1):1–73, July 1929. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
1.1;   http://rmp.aps.org/abstract/RMP/v1/i1/p1_1. See er-

rata [Bir29a].

Compton:1929:CPL

[Com29] Arthur H. Compton. The corpuscular properties of light. Reviews of Modern Physics, 1(1):74–89, July 1929. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.1.74;
 http://rmp.aps.org/abstract/RMP/v1/i1/p74_1.

Darrow:1929:STM

[Dar29] Karl K. Darrow. Statistical theories of matter, radiation and electricity. Reviews of Modern Physics, 1(1):90–155, July 1929. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.1.90;
 http://rmp.aps.org/abstract/RMP/v1/i1/p90_1.

Kemble:1929:GPQ

[Kem29] Edwin C. Kemble. The general principles of quantum mechanics. Part I. Reviews of Modern Physics, 1(2):157–215, October 1929. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
1.157;   http://rmp.aps.org/abstract/RMP/v1/i2/p157_1. See

errata [Kem30].

Kemble:1930:EPG

[Kem30] Edwin C. Kemble. Errata to Part I, general principles of quantum mechanics. Reviews of Modern Physics, 2(1):59, January 1930. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.2.59;
 http://rmp.aps.org/abstract/RMP/v2/i1/p59_1. See [Kem29].

Mohler:1929:RPI

[Moh29] F. L. Mohler. Recombination and photo-ionization. Reviews of Modern Physics, 1(2):216–227, October 1929. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.1.
216;   http://rmp.aps.org/abstract/RMP/v1/i2/p216_1.

Osgood:1929:SSX

[Osg29] Thomas H. Osgood. Spectroscopy of soft X-rays. Reviews of Modern Physics, 1(2):228–240, October 1929. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.1.228;   http://rmp.
aps.org/abstract/RMP/v1/i2/p228_1.